It took Canoe A 1.5hrs to travel from Duck Wharf to Buglers Bay. Its rate in still water was 12km/hr but it was going with the current. Canoe B was traveling at 14 km /hr in still water but it was going from Burglers Bay to Duck Wharf and the current was going against it. It took Canoe B 5 hrs to make the trip. What was the rate of the current?
thanx

Answer by Paul(988)

Let x be the speed of the current:
EQUATION:
1.5(12+x)=5(14-x)
18+1.5x=70-5x
6.5x=52
x=8
Hence, the speed of the current was 8mph.
Paul.